The high temperatures (in degrees Fahrenheit) of a randon sample of 10 small towns are: 96.7 98.6 97.3 96.8 97.6 99.8 99.9 96.9 99.1 98.9 Assume high temperatures are normally distributed. Based on this data, find the 80% confidence interval of the mean high temperature of towns Enter your answer as an open-interval (ie, parentheses) acurate to two decimal places (because the sample data are reported accurate to one decimal place). 80% c.1. Answer should be obtained without any preliminary rounding, However, the critical value may be rounded to 3 decimal places Poires possible: 1 This is attempt 1 of 3. Poss shis question to fonum
Answer #1

Solution:- Given that values 96.7,98.6,97.3,96.8,97.6,99.8,99.9,96.9,99.1,98.9

x = 98.16 , S = 1.2456 , n = 10

df = n-1 = 10-1 = 9

t = 1.383

80% confidence interval of the mean high temperature : x +/- t*s/sqrt(n)
: 98.16 +/- 1.383*1.2456/sqrt(10)
: 97.615 , 98.705